ELECTRONIC SPECTROSCOPY OF MOLYBDENUM MONOCARBIDE

\underline{DALE J. BRUGH}, National Institute of Standards and Technology, Optical Technology Division, Gaithersburg, MD 20899; TJ RONNINGEN and MICHAEL D. MORSE, University of Utah, Department of Chemistry, Salt Lake City, UT 84112.

In a continuing effort to understand the nature of the metal carbon bond in transition metal monocarbides we have undertaken the first spectroscopic investigation of MoC by resonant two-photon ionization spectroscopy. Molybdenum carbide was produced by laser vaporization of a Mo sample disk in a supersonic expansion of He and 3% CH_{4}. Using 6.42 eV photons for photoionization we have observed a number of transitions between 17 700 cm^{−1} and 24 000 cm^{−1}. Twenty four of the observed bands have been studied at sufficient resolution (0.04 cm^{−1}) to allow the rotational structure of each transition to be resolved. An analysis of the isotope shifts has allowed a tentative identification of three band systems among the 24 observed transitions. A number of the observed bands have remained ungrouped. Every rotationally resolved band appears to be an Ω = 1 ← 0 transition. Given that the ground state of NbC^{a} is known to be 11σ^{2}2δ^{1}, 2Δ_{r} and that of RuC^{b} is known to be 11σ^{2}2δ^{4}, ^{1}Σ^{+}, it seems likely that the addition of another electron on moving from NbC to MoC yields for a ground state the Ω = 0^{+} component of a ^{3}Σ^{−} term. This is consistent with the observation that every transition in this study originates from an Ω = 0 state. The X^{3}Σ^{−} rotational constant for the most abundant isotope, ^{98}Mo^{12}C, has been determined to be 0.553640 ± 0.000055 cm^{−1}. This corresponds to a ground state bond length of 1.687719 ± 0.000084 Å.

aB. Simard, P. I. Presunka, H. P. Loock, A. Berces, and O. Launila, submitted to J. Chem. Phys. (1997).

bJ. Langenberg, R. S. Dabell, L. Shao, D. Dreessen, and M. D. Morse, in preparation.